# This work has been submitted to the IEEE Access for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible. xx.xxxx/ACCESS.2024.xxxxxxx

### v2

@@ I-C Paper Structure
The rest of this survey paper is organized as shown in Fig. 3. Section II delves into the CL fundamentals. Section III provides an overview of SNNs and the state-of-the-art NCL methods. In Section IV, a summary of real-world application use-cases of NCL is presented. Then, Section V discusses open research challenges for meeting the CL desiderata with SNNs. Finally, Section VI presents the conclusion.
@@ II Continual Learning
In recent years, DNNs have revolutionized the field of AI [55, 51, 56], due to their effective training technique (i.e., gradient descent-based backpropagation), while assuming the training data are independently and identically distributed (i.i.d) [57, 58]. However, a critical question arises: how can we integrate new data into existing models? One approach is to retrain with new data, while using current parameters as the initial state [58]. However, successive training tends to cause NNs to suffer from CF [59, 60]. Although retraining from scratch effectively tackles CF issues, but it is inefficient and almost impossible in some cases, such as real-time learning scenarios where the model must update and adapt to continuous data streams instantly with limited resources (e.g., developmental learning for autonomous agents [61]).
@@ II-B Desiderata of CL
In this section, we explain the desired characteristics (desiderata) of a CL algorithm, and show an overview of key requirements to achieve CL capability using a robotic arm as an example in Fig. 4.
@@ II-B2 No/Minimal Usage of Old Data
Minimizing or eliminating the reliance on data from previous tasks is crucial for a CL algorithm, because of storage constraints and privacy concerns. Recent researches have made progress to achieve this goal [68, 69, 70, 71, 72, 73]. For instance, previous studies utilized mutual information maximization and compressed gradients to minimize the reliance on old data [68, 69, 70], and employed representative subsets to preserve knowledge [71].
@@ II-B4 Positive Forward Transfer (FWT)
$\text{FWT}_{k}=\frac{1}{k-1}\sum_{j=2}^{k}(a_{j,j}-\tilde{a}_{j})$ (1)
where, $\tilde{a}_{j}$ is the accuracy of a randomly-initialized model trained with data $\mathcal{D}_{j}$ for the $j$-th task [16]. Recent researches explored different methods to enable FWT, including studies in [82, 83, 84, 85, 86, 87, 88, 53, 89, 90, 91].
@@ II-B5 Positive Backward Transfer (BWT)
A CL algorithm should transfer knowledge from later tasks to past tasks for improving performance on the past tasks, known as Positive Backward Transfer (BWT). Here, zero backward transfer suggests zero forgetting [53]. Formally, $\text{BWT}_{k}$ evaluates the average influence of learning the $k$-th task on all old tasks, and can be stated as:
@@
where, $a_{k,j}$ is the accuracy of a model at $j$-th task after incremental learning of the $k$-th task [16].
@@ II-B6 Controlled Forgetting
A CL algorithm should forget old yet insignificant knowledge to make room for learning new information, known as controlled forgetting. Although humans do not experience sudden memory loss, a gradual decrease in memory (forgetting) over time is natural, hence controlled forgetting is beneficial for managing memory retention [92]. Related methods include [93, 94, 95] for DNNs and [96, 97, 41, 42] for SNNs.
@@ II-B7 Fast Adaptation and Recovery
II-C Scenarios in CL
CL scenarios can be categorized using a taxonomy presented in Fig. 5, while their formal comparison is provided in Table II. The initially-identified scenarios are Task-Incremental Learning (Task-IL), Domain-Incremental Learning (Domain-IL), and Class-Incremental Learning (Class-IL). While these scenarios cover some desiderata of CL, they fail to meet the requirement of TA [53]; see Fig. 6. They assume that tasks have clear and well-defined boundaries during training. Therefore, recent studies proposed expanding the CL scenarios by introducing Discrete TA (DTA) [74], Continuous TA (CTA) [74], and Open-World TA (OWTA) [53].
@@ II-C Scenarios in CL
II-C1 Task-Incremental Learning (Task-IL)
@@
II-C2 Domain-Incremental Learning (Domain-IL)
@@
II-C3 Class-Incremental Learning (Class-IL)
@@
II-C4 Discrete Task Agnostic (DTA)
@@
II-C5 Open-World Task Agnostic (OWTA)
@@
II-C6 Continuous Task Agnostic (CTA)
@@
II-D Online Continual Learning (OCL)
In OCL, a model learns from a continuous data stream, so that each input sample during inference is used as training data for updating the models’ knowledge [103, 104, 46]. As a new sample arrives, it is immediately used for inference, and for updating the model on-the-fly. This capability is highly desired for systems that face highly dynamic environments (e.g., autonomous mobile agents), where new data is encountered in real-time and immediate updates to the model are required. Its challenges include integrating new knowledge and retaining previous information, while considering limited memory and power/energy budgets. In this survey, we emphasize that OCL is the expected capability for autonomous embedded systems.
II-E CL Methods
In this section, we discuss existing methods for addressing CF in various CL scenarios. These methods are categorized into five approaches based on representation, regularization, rehearsal/replay, optimization, and architecture [16], as follows.
II-E1 Representation-based Approach
@@
    Self-Supervised Learning (SSL): It makes the model learn to generate useful data representations without relying on explicit labels, by creating supervisory signals from the input data to uncover the underlying structure of the data [105, 106]. Related studies are discussed in [107, 108, 109, 110, 111, 112, 113, 114, 115].
@@
    Pre-training for Downstream CL: It is the initial phase of model training on a large diverse dataset to learn general-purpose representations, which can be fine-tuned for specific downstream tasks, hence having strong knowledge transfer [83, 116, 117, 118]. To maintain generalizability for future tasks, several strategies have been developed, such as Pre-trained Representations [119, 120, 88, 121, 122], Task-adaptive Prompts [16, 123, 124, 125, 126, 127, 128], Saving Prototypes and Enhancing Classifiers [129, 130, 131], and Optimizing an Updatable Backbone [132, 133, 83, 134, 135].
@@
II-E2 Regularization-based Approach
@@
    $\theta_{t}=\arg\min_{\theta}\left(\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\mathcal{L}_{t}(f_{\theta}(x_{i}),y_{i})+\lambda\mathcal{R}_{f}(f_{\theta},\mathcal{D}_{1:t-1})\right)$ (3)
    where, $\theta_{t}$ are the model parameters after learning task $t$, $\mathcal{L}_{t}$ is the loss function for task $t$, $N_{t}$ is the number of samples in task $t$, $\mathcal{R}_{f}(f_{\theta},\mathcal{D}_{1:t-1})$ is the function regularization term that penalizes changes in the models’ function based on its performance on previous tasks $\mathcal{D}_{1:t-1}$ and $\lambda$ is a hyperparameter controlling the strength of the regularization.
@@
    Besides the above prominent methods, several alternatives have also been proposed: Sequential Bayesian Inference over Function Space [16, 161, 162, 163] and Conditional Generation [164, 165, 166].
II-E3 Rehearsal/Replay-based Approach
@@
    Experience Replay (ER): It stores a few old training samples in a small memory buffer. It includes several strategies such as Sample Selection (SS) [85, 167, 86, 157], Gradient/Optimization-based Techniques [16, 103, 168, 169, 170, 171, 172], Storage Efficiency [173, 174, 174, 65, 171, 175, 176], Integration of ER with KD [157, 177], Mitigation of Data Imbalance [178, 179, 180, 180, 165, 181, 112, 182, 183, 184, 185], Learning Plasticity Enhancement [186, 187], and Overfitting Alleviation [188, 189, 190, 191].
@@
    Generative Replay: It uses an additional generative model such as Generative Adversarial Networks (GANs) [192] or Variational Autoencoders (VAEs) [193] to replay the previously-learned data [21]. There are several GANs-based techniques as studied in [194, 164, 64, 195, 196]. However, they usually suffer from label inconsistency [64, 197]. Toward this, other techniques employ autoencoder-based strategies [198, 197, 21].
@@
II-E4 Optimization-based Approach
@@
II-E5 Architecture-based Approach
@@
All these existing methods for fulfilling the desiderata of CL are also summarized and presented in Table III.
TABLE III: Summary of CL methods based on Representation-, Regularization-, Rehearsal/Replay- and Architecture-based approaches that potentially fulfill “+” or have fulfilled “✓” the desiderata of CL.
@@
II-F CL Settings
II-F1 Supervised Learning
II-F2 Unsupervised Learning
II-F3 Reinforcement Learning (RL)
II-G Evaluation Metrics
To evaluate the performance of an efficient CL system, metrics are essential. To assess the overall performance of a CL algorithm, average accuracy (AA) [264, 86] and average incremental accuracy (AIA) [157, 178] are often used. Model size is used to measure the memory size for storing the model parameters [265], while sample storage size (SSS) is used to measure the storage size for storing data samples. To evaluate the computational resources required for the training and inference phases, metrics such as the number of floating-point operations-per-second (FLOPS) and time complexity are often used. Meanwhile, power/energy efficiency measures the efficiency gains of a CL system during its operational life-time.

### v3

@@ I-C Paper Structure
The rest of this survey paper is organized as shown in Fig. 3. Section II delves into the CL background. Section III provides detailed technical background of neuromorphic computing systems across multiple aspects and comprehensively reviews state-of-the-art NCL methods, with comparative quantitative and trade-off analysis. In Section IV, several application use-cases of NCL are provided and analyzed. Then, Section V discusses open challenges for meeting the CL desiderata with SNNs and proposes future research directions for each challenge. Finally, Section VI presents the conclusion.
@@ II Continual Learning
In recent years, ANNs have revolutionized the field of AI [8, 51, 61], due to their effective training technique (i.e., gradient descent-based backpropagation), while assuming the training data are independently and identically distributed (i.i.d) [62, 63]. However, a critical question arises: how can we integrate new data into existing models? One approach is to retrain with new data, while using current parameters as the initial state [63]. However, successive training tends to cause NNs to suffer from CF [64, 65]. Although retraining from scratch effectively tackles CF issues, but it is inefficient and almost impossible in some cases, such as real-time learning scenarios where the model must update and adapt to continuous data streams instantly with limited resources (e.g., developmental learning for autonomous agents [66]).
@@ II-B Desiderata of CL
In this sub-section, we address the key question Q1 by explaining the desired characteristics (desiderata) of a CL algorithm, and show an overview of key requirements to achieve CL capability using a robotic arm as an example in Fig. 4.
@@ II-B2 No/Minimal Usage of Old Data
Minimizing or eliminating the reliance on data from previous tasks is crucial for a CL algorithm, because of storage constraints and privacy concerns. Recent research has made progress to achieve this goal [73, 74, 75, 76, 77, 78]. For instance, previous studies used mutual information maximization and compressed gradients to minimize the reliance on old data [73, 74, 75], and employed representative subsets to preserve knowledge [76].
@@ II-B4 Positive Forward Transfer (FWT)
$\text{FWT}_{k}=\frac{1}{k-1}\sum_{j=2}^{k}(a_{j-1,j}-\tilde{a}_{j})$ (1)
where, $\tilde{a}_{j}$ is the accuracy of a randomly-initialized model trained with data $\mathcal{D}_{j}$ for the $j$-th task [16]. The parameter $a_{j-1,j}$ denotes accuracy on task $j$ after learning task $j-1$, but before task $j$ and $\tilde{a}_{j}$ denotes baseline performance on task $j$ without prior learning. High FWT indicates that the model generalizes well to unseen tasks by reusing learned knowledge. Recent research explored different methods to enable FWT, including studies in [87, 88, 89, 90, 91, 92, 93, 53, 94, 95, 96].
@@ II-B5 Positive Backward Transfer (BWT)
A CL algorithm should transfer knowledge from later tasks to past tasks for improving performance on the past tasks, known as Positive Backward Transfer (BWT). Negative BWT reflects CF and zero BWT suggests zero forgetting [53]. Formally, $\text{BWT}_{k}$ evaluates the average influence of learning the $k$-th task on all old tasks, and can be stated as:
@@
where, $a_{k,j}$ is the accuracy of a model at $j$-th task after incremental learning of the $k$-th task [16]. The parameter $a_{j,j}$ denotes accuracy of task $j$ when it was first learned.
@@ II-B6 Controlled Forgetting
A CL algorithm should forget old yet insignificant knowledge to make room for learning new information, known as controlled forgetting. Although humans do not experience sudden memory loss, a gradual decrease in memory (forgetting) over time is natural, hence controlled forgetting is beneficial for managing memory retention [97]. Related methods include [98, 99, 100] for ANNs and [101, 102, 42, 43] for SNNs.
@@ II-B7 Fast Adaptation and Recovery
II-C CL Settings
In this sub-section, we address the key question Q1 by introducing the definition of learning settings in which the CL methods operate, based on how the data label and information are utilized for learning process, including the supervised, unsupervised, and reinforcement learning settings.
II-C1 Supervised Learning
II-C2 Unsupervised Learning
II-C3 Reinforcement Learning (RL)
II-D Evaluation Metrics
To evaluate the performance of an efficient CL system, metrics are essential. This sub-section addresses the key question Q1 for evaluation metrics. To assess the overall performance of a CL algorithm, average accuracy (AA) [107, 91], average incremental accuracy (AIA) [108, 109] and forgetting (F) are often used, where F is the average maximum drop in classification accuracy over time, with lower values indicating better retention. BWT and FWT are also used to quantify memory stability and learning plasticity, respectively, in CL. Model size is used to measure the memory size for storing the model parameters [110], while sample storage size (SSS) is used to measure the storage size for storing data samples. To evaluate the computational resources required for the training and inference phases, metrics such as the number of floating-point operations-per-second (FLOPS) and time complexity are often used. Meanwhile, power/energy efficiency measures the efficiency gains of a CL system during its operational life-time.
II-E Scenarios in CL
In this sub-section, we address key question Q1 regarding the CL scenarios by presenting their definitions, taxonomy (shown in Fig. 5), and formal comparison in Table IV. The initially-identified scenarios are Task-Incremental Learning (Task-IL), Domain-Incremental Learning (Domain-IL), and Class-Incremental Learning (Class-IL). While these scenarios cover some desiderata of CL, they fail to meet the requirement of TA [53]; see Fig. 6. They assume that tasks have clear and well-defined boundaries during training. Therefore, recent studies proposed expanding the CL scenarios by introducing Discrete TA (DTA) [79], Continuous TA (CTA) [79], and Open-World TA (OWTA) [53].
@@ II-E Scenarios in CL
II-E1 Task-Incremental Learning (Task-IL)
@@
II-E2 Domain-Incremental Learning (Domain-IL)
@@
II-E3 Class-Incremental Learning (Class-IL)
@@
II-E4 Discrete Task Agnostic (DTA)
@@
II-E5 Open-World Task Agnostic (OWTA)
@@
II-E6 Continuous Task Agnostic (CTA)
@@
II-F Online Continual Learning (OCL)
In OCL, a model learns from a continuous data stream, so that each input sample during inference is used as training data for updating the models’ knowledge [115, 116, 46]. As a new sample arrives, it is immediately used for inference, and for updating the model on-the-fly. This capability is highly desired for systems that face highly dynamic environments (e.g., autonomous mobile agents), where new data is encountered in real-time and immediate updates to the model are required. Its challenges include integrating new knowledge and retaining previous information, while considering limited memory and power/energy budgets. In this survey, we emphasize that OCL is the expected capability for embedded AI systems.
II-G CL Methods
In this sub-section, we address the key question Q2, by providing an overview of the existing methods that address CF in various CL scenarios, and the desiderata they fulfill; summarized in Table V. These methods are categorized into five approaches based on representation, regularization, rehearsal/replay, optimization, and architecture [16], as follows.
II-G1 Representation-based Approach
@@
    Self-Supervised Learning (SSL): It makes the model learn to generate useful data representations without relying on explicit labels, by creating supervisory signals from the input data to uncover the underlying structure of the data [117, 118]. Related studies are discussed in [119, 120, 121, 122, 123, 124, 125].
@@
    Pre-training for Downstream CL: It is the initial phase of model training on a large diverse dataset to learn general-purpose representations, which can be fine-tuned for specific downstream tasks, hence having strong knowledge transfer [88, 126, 127, 128]. To maintain generalizability for future tasks, several strategies have been developed, such as Pre-trained Representations [129, 130, 93, 131, 132], Task-adaptive Prompts [133, 134, 135, 136, 137, 138], Saving Prototypes and Enhancing Classifiers [139, 140, 141], and Optimizing an Updatable Backbone [142, 143, 88, 144, 145].
@@
II-G2 Regularization-based Approach
@@
    $\bm{\theta}_{t}=\arg\min_{\bm{\theta}}\left(\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\mathcal{L}_{t}\big{(}f_{\bm{\theta}}(\mathbf{x}_{i}),y_{i}\big{)}+\lambda\,\mathcal{R}_{f}\big{(}f_{\bm{\theta}},\mathcal{D}_{1:t-1}\big{)}\right)$ (3)
    where, $\bm{\theta}\in\mathbb{R}^{d}$ denotes the model parameters (weights) after learning task $t$. The term $\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\mathcal{L}_{t}(f_{\theta}(\mathbf{x}_{i}),y_{i})$ calculates the average loss on the current task $t$, where $N_{t}$ is the number of samples in task $t$, $\mathbf{x}_{i}\in\mathbb{R}^{n}$ is the input feature vector for the $i^{\text{th}}$ training sample for task $t$, $y_{i}$ is the corresponding ground truth label for $\mathbf{x}_{i}$, $f_{\theta}(\mathbf{x}_{i})$ is the model’s prediction for input $\mathbf{x}_{i}$ using parameters $\bm{\theta}$ and $\mathcal{L}_{t}$ is the task-specific loss function (e.g., cross-entropy for classification). $\mathcal{R}_{f}(f_{\theta},\mathcal{D}_{1:t-1})$ is the function regularization term that penalizes changes in the models’ function $f_{\theta}$ based on its performance on previous data $\mathcal{D}_{1:t-1}$ from tasks 1 to $t-1$ and $\lambda$ is a hyperparameter controlling the strength of the regularization i.e., the balance between stability and plasticity. The goal is to find the set of weights $\bm{\theta}_{t}$ that minimize both the loss on the current task and the deviation from prior knowledge.
@@
    Besides the above prominent methods, several alternatives have also been proposed: Sequential Bayesian Inference over Function Space [169, 170, 171] and Conditional Generation [172, 173, 174].
@@ II-G3 Rehearsal/Replay-based Approach
    Experience Replay (ER): It stores a few old training samples in a small memory buffer. It includes several strategies such as Sample Selection (SS) [90, 175, 91, 108], Gradient/Optimization-based Techniques [115, 176, 177, 178, 179, 180], Storage Efficiency [181, 182, 70, 179, 183, 184], Integration of ER with KD [108, 185], Mitigation of Data Imbalance [109, 186, 187, 173, 188, 124, 189, 190, 191, 192], Learning Plasticity Enhancement [193, 194], and Overfitting Alleviation [195, 196, 197, 198, 199].
@@
    Generative Replay: It uses an additional generative model such as Generative Adversarial Networks (GANs) [200] or Variational Autoencoders (VAEs) [201] to replay the previously-learned data [21]. There are several GANs-based techniques as studied in [202, 172, 69, 203, 131]. However, they usually suffer from label inconsistency [69]. Toward this, other techniques employ autoencoder-based strategies [204, 205, 21].
@@
II-G4 Optimization-based Approach
@@
II-G5 Architecture-based Approach
@@
II-H Limitations of Existing CL Methods
The CL desiderata fulfilled by the existing CL methods (discussed in Section II-G) are presented in Table V. This table shows that, different CL methods feature different advantages and disadvantages. Furthermore, the limitations of existing CL methods in DNNs and the challenges that impede their direct application to SNNs are discussed as follows, answering key question Q3. A summary of these limitations is provided in Table VI. For example, weight regularization with methods like EWC [72] and SI [68] requires estimating and storing importance measures (e.g., FIM calculations) for every synapse, adding a quadratic penalty term for each learned task, which is resource-intensive in high-dimensional, event-driven SNNs, therefore has a linearly increasing computational cost [44]. Moreover, such importance scores assume a static activation landscape, which is unstable in SNNs due to variable spiking activity and dynamic thresholds. Replay-based methods (e.g., iCaRL [108], DGR [202], GEM [91]) require storing and reprocessing raw input samples or features. This is not well suited for neuromorphic hardware, which operates under strict memory and power constraints and lacks the infrastructure to store dense data or continuously replay high-dimensional inputs efficiently. Many CL methods in DNNs rely on precise gradient computations through differentiable activation functions. In contrast, SNNs involve discontinuous spike events and temporal dynamics, making conventional BP infeasible or biologically implausible. Although SG methods exist, they are computationally expensive. Moreover, spike-based learning involves temporally extended input encoding. Mapping gradient-based loss functions from DNN-CL to temporally sparse spike trains introduces ambiguity in assigning credit across time, which is non-trivial and often requires costly unrolling (e.g., through BPTT). Most DNN-based CL methods assume synchronous, frame-based computation. SNNs operate asynchronously, responding to sparse events. This fundamental mismatch makes it difficult to translate CL mechanisms designed for batch updates and full forward/backward passes into the SNN regime without re-engineering their entire workflow.
These limitations in adapting conventional CL strategies to SNNs underscore the critical need for a paradigm shift towards neuromorphic-specific CL algorithms that can achieve higher performance while minimizing energy consumption. This is especially important for deployments in embedded AI systems (e.g., mobile robots/agents and IoT-Edge) where compute and memory budgets are limited, highlighting the reasons why the energy-efficient CL algorithms are crucial for resource-constrained embedded AI systems posed by the key question Q3.
Motivated by the limitations of existing CL methods, NCL concept has emerged as a potential solution as it has characteristics that align with the desiderata of CL. Specifically, NCL leverages SNN computation models which has event-driven operations, thus enabling energy-efficient learning. Furthermore, SNNs can perform unsupervised learning due to their bio-plausible learning rules. Details of NCL concept and review of NCL methods that are explicitly designed to operate under the constraints of event-driven computation, limited memory, and hardware variability inherent to neuromorphic systems will be discussed in Section III.
TABLE V: Representation-, Regularization-, Rehearsal/Replay- and Architecture-based CL methods that potentially fulfill “+” or have fulfilled “✓” the desiderata of CL.
@@ II-H Limitations of Existing CL Methods
TABLE VI: Summary of ANN-based CL methods, outlining their computational/memory overhead.
Approach CL Method Network/Model Limitations
Regularization-based EWC [72],
SI [68], MAS [153],
RWalk [107],
R-EWC [156],
XK-FAC [157],
ALASSO
 [158],
GD [168]
MLP, CNN
LeNet, VGG-16
ResNet-18
ResNet-50
ResNet-18
Wide Residual Network
- Linearly increasing computations with the number of parameters and tasks due to Fisher Information Matrix (FIM) calculations.
- High computations due to factorized rotation of the parameter space to diagonalize the FIM.
- High computations due to the complexity of handling batch normalization.
- High computations due to more accurate quadratic approximation of the loss function for improving performance.
- High computations due to large-scale distillation over unlabeled data.
Replay-based GEM [91],
A-GEM[222],
DGR [202],
RAR [183],
DGM [69],
L-VAEGAN [21],
EEC [205]
MLP, ResNet-18
GAN,
MLP, ResNet-18
DCGAN, ResNet-18,
VAE-GAN,
DCGAN, ResNet-18
- High computations and memory overhead during training due to the requirement for backward passes.
- High computations due to adversarial training, high memory for storing and generating past data.
- High computations due to adversarial training with sparse attention masks.
- High computations due to combination of VAEs with GANs.
- High computations and memory overhead due to encoding and decoding of entire past episodes.
Representation-based Co2L [124],
DualNet [119]
ResNet-18,
Two parallel CNNs
- High computations due to contrastive learning.
- Memory overhead due to maintaining two networks.
TwF [93],
GAN-Memory [131], CODA-Prompt [134], S-Prompts [136], Barlow Twins + EWC [146]
ResNet-18,
GAN,
Vision Transformer (ViT),
Transformer-based
ResNet-50
- Memory overhead due to parallel networks.
- High computations due to generative modeling.
- High computations due to attention calculations of ViTs leading to quadratic complexity in terms of input size.
- High memory overhead as prompts grow with tasks.
- High computations due to quadratic penalty in EWC.
Architecture-based PNN [94],
PathNet [95],
Model Zoo [246]
MLP, CNN
CNN,
Wide-Resnet
- Exponential memory increase due to new sub-networks added for new tasks.
- High computations due to multiple parallel networks.
- Linearly increasing memory due to a separate sub-network storage for each task, and ensemble of all sub-networks.